OpenStudy (anonymous):

The rectangular coordinates of a point are given. Find polar coordinates for the point. 1. (4,0) 2.(3,4) 3.(2,-2)

OpenStudy (anonymous):

Rectangular to polar: \[r=\sqrt{x^2+y^2}\] \[\theta=\tan^{-1} \frac{ y }{ x }\] Polar to rectangular \[x=r \cos \theta\] \[y = r \sin \theta\]

OpenStudy (anonymous):

I think I got it. Is 1. (4,0) 2. (5,53.1) 3. sqrt8,45) ?

OpenStudy (anonymous):

1 and 2 are good. For 3, (2, -2) is in the 4th quadrant, so the angle would be -45° or one of its coterminal angles.

OpenStudy (anonymous):

alright thanks:)

OpenStudy (anonymous):

you're welcome